OpenStudy (anonymous):

The numerator and denominator of a fraction are in the ratio of 3 to 5. If the numerator and denominator are both increased by 2, the fraction is now equal to 2/3. If n = the numerator and d = the denominator, which of the following systems of equations could be used to solve the problem? a) 5n = 3d and 3n + 6 = 2d + 4 b) 5n = 3d and 4n + 4 = 3d + 6 c) 3n = 5d and 3n + 6 = 2d + 4

OpenStudy (anonymous):

"If the numerator and denominator are both increased by 2, the fraction is now equal to ." Now equal to what?

OpenStudy (anonymous):

Oops. :$ The fraction never copied over. It's 2/3.

OpenStudy (anonymous):

Some fraction \[\frac{ a }{ b }\ = \frac{ 3 }{ 5 }\] add 2 to the top and bottom\[\frac{ a+2 }{ b+2 }\ = \frac{ 2 }{ 3 }\] Cross multiply each fraction so there aren't any more denominators.